### v1

# Knowledge-Centric Templatic Views of Documents

Abstract
Authors seeking to communicate with broader audiences often compose their ideas about the same underlying knowledge in different documents and formats – for example, as slide decks, newsletters, reports, brochures, etc. Prior work in document generation has generally considered the creation of each separate format to be different a task, developing independent methods for generation and evaluation. This approach is suboptimal for the advancement of AI-supported content authoring from both research and application perspectives because it leads to fragmented learning processes, redundancy in models and methods, and disjointed evaluation. Thus, in our work, we consider each of these documents to be templatic views of the same underlying knowledge, and we aim to unify the generation and evaluation of these templatic views of documents. We begin by introducing an LLM-powered method to extract the most important information from an input document and represent this information in a structured format. We show that this unified representation can be used to generate multiple templatic views with no supervision and with very little guidance, improving over strong baselines. We additionally introduce a unified evaluation method that is template agnostic, and can be adapted to building document generators for heterogeneous downstream applications. Finally, we conduct a human evaluation, which shows that humans prefer 82% of the downstream documents generated with our method. Furthermore, the newly proposed evaluation metric correlates more highly with human judgement than prior metrics, while providing a unified evaluation method.
Figure 1: Visualization of our method. Given an input document, we prompt the LLM to generate a Structure Unified Representation (SURe). We can use the SURe to prompt the model to generate a templatic view of the input document. We then evaluate the generations using our unified evaluation.
1 Introduction
Sharing information in multiple formats is an important part of communication in a variety of domains, as it allows for knowledge to be disseminated to a wider audience. A product manager may need to create a requirements document, a product pitch deck, and a product announcement newsletter. Likewise, a person on the job market may create a resume, a cover letter, and a personal website. We consider these documents to be templatic views of the same underlying knowledge.
This is also true for the scientific domain, in which researchers need to create documents in multiple formats to effectively disseminate their work, – such as through conference talks, social media posts, poster presentations, and non-technical blog posts. Sharing knowledge in multiple formats broadens its audience and can help bridge the information gap between domain experts, researchers in adjacent fields, and even the general public – leading to greater understanding, new collaboration and accelerated progress in the field. With the significant increase in the pace of publication in recent years Bornmann and Mutz (2014), it has become essential for scientists to engage with the community and share their work in more compact, efficient, or easier to digest formats.
Past work on document generation has focused on developing methods of generation and evaluation for each specific document type individually Fu et al. (2021); Qiang et al. (2016); Chandrasekaran et al. (2020). However, tailoring methods to each template is time consuming and limiting, resulting in methods that are usable only for each intended template that cannot be compared across template types with specialized evaluation metrics. Our goal is to unify the tasks of generating and evaluating templatic views of documents. By doing so, we aim to allow system designers to adapt to new document types, templates and domains easily and efficiently, and to allow researchers to compare approaches more flexibly and fairly.
We begin by setting a few desiderata for our system for templatic view generation; it must be (1) unified, (2) adaptable, and (3) require minimal supervision. In order to design such a system to generate each of these templates, the method must be able to create structure and follow style norms associated with each template. We thus introduce an LLM-powered approach to generate a Structure Unified Representation (SURe) that represents the most important knowledge of an input document in a structured format. We demonstrate that this SURe can be used with no supervision and with very little prompted guidance about the output formats, to generate multiple templatic views of a document.
Meanwhile, automatic evaluation is crucial for researchers and developers to iterate more quickly than they otherwise would be capable of by relying solely on human evaluation. However, because in the past each templatic view has been tackled independently, previous work has used different evaluation metrics for each task, such as BERTScore for summarization Zhang* et al. (2020), Slide Level ROUGE for slide decks Qiang et al. (2016), or even a custom-trained, model-based metric for posters Wang et al. (2015). In this paper, we introduce Template Agnostic Evaluation (TAE) – a unified precision-recall style method of automatic evaluation that is template agnostic. Importantly, this evaluation framework is highly customizable, allowing users to easily integrate existing text-based metrics from the literature into its formulation and tailor its function to their specific use case.
We evaluate our unified approach to templatic view generation and evaluation on 3 types of documents: slides, posters, and blog posts. In each of these, the inputs are scientific articles and the outputs are expected to be documents of the corresponding types. We use the DOC2PPT Fu et al. (2021), Paper-Poster Qiang et al. (2016), and LongSumm Chandrasekaran et al. (2020) datasets, respectively. We show that our method of generating a Structured Unified Representation leads to improvements in performance across each tasks, demonstrating that LLMs are capable of generating templatic views with structure-augmented prompts, despite having no supervision. We additionally conduct a human evaluation to validate both our document generation method and evaluation metric. We show that humans prefer the output yielded by our method 82% of the time and that our evaluation metric correlates more highly with human preference than other popular metrics, such as ROUGE and BERTScore Lin (2004); Zhang* et al. (2020).
While the focus of this paper is the transformation of scientific documents, this choice was primarily driven by the availability of publicly available data. We note that nothing about our unified framework is domain specific, and we believe our approach for the creation and evaluation of templatic views of documents is easily adaptable to novel domains, and document types. More generally, we believe our work is a step towards the broader and more challenging goal of AI-assisted authoring support, and future research in reasoning over and generation of long documents.
2 Related Work
There are several areas of related research in NLP that are relevant to the problems of document transformation and evaluation.
Document summarization has been explored in a number of domains, including news See et al. (2017), literature Sciré et al. (2023), law Deroy et al. (2023), and dialogue Chen et al. (2021). In the scientific domain, summarization of scientific papers has taken the form of long form summaries Chandrasekaran et al. (2020), abstract generation Cohan and Goharian (2015), conference talks Lev et al. (2019), and query based summaries Fok et al. (2023). These summaries can be either extractive Sefid and Giles (2022) or abstractive Chandrasekaran et al. (2020).
Although the tasks of slide and poster generation have generally been considered separate from scientific summarization, they are related in that both tasks require taking an input article, then organizing and abstracting the information to generate a new document. Past work has developed methods for slide generation from papers Hu and Wan (2015); Li et al. (2021); Hu and Wan (2015); Fu et al. (2021), from code Wang et al. (2023a), or based on a query Sun et al. (2021). Poster generation has been explored in the form of content extraction for posters Xu and Wan (2021), interactive generation Wang et al. (2015), or full content generation using graphical models Qiang et al. (2016). To the best of our knowledge, our work is the first to create a unified method capable of generating a diverse range of templatic views of a source document.
Large Language Models (LLMs), which are central to our approach, have shown impressive capabilities in a variety of tasks Radford et al. (2019); Brown et al. (2020). Based on the transformer architecture Vaswani et al. (2017), LLMs have shown emergent abilities in tasks such as arithmetic and question answering Wei et al. (2022a). Similar to chain of thought prompting Wei et al. (2022b) and content planning prompting Wang et al. (2023b), we show that by generating an intermediate Structured Unified Representation (SURe) of an input document can improve performance over simply prompting the model to generate the final document Additionally, past work has explored using knowledge graphs to provide structure to LLMs in order to improve performance Pan et al. (2024); Rosset et al. (2020); Lewis et al. (2020). Similarly, we show that the structure provided by the SURe improves the quality of the final document, while being more easily adapted to new domains than a knowledge graph.
As past work has considered generation of templatic views to be separate tasks, methods for automatic evaluation of different templatic views has also been considered independently. LongSumm, the shared task introduced by Chandrasekaran et al. (2020), uses ROUGE scores to evaluate model performance Lin (2004). Fu et al. (2021) introduced Slide Level ROUGE to evaluate slide generation, which is a variant of ROUGE that contains a penalty for the number of slides. Qiang et al. (2016) used a trained regressor to predict the size of panels in a poster. In the summarization space, many automatic metrics for evaluation have been introduced such as BERTScore Zhang* et al. (2020), UniEval Zhong et al. (2022), BARTScore Yuan et al. (2021), BLANC Vasilyev et al. (2020), and MoverScore Zhao et al. (2019). However, these metrics are intended for a simple input document-summary setup, and do not take into account factors that affect the quality of other types of documents, such as structure and ordering. Our work is the first to introduce template agnostic evaluation, allowing uniform comparison of performance across template types.
Prompt Function Prompt
Generate the SURe
"Given the input text, extract the document title and authors.
For each section in the given input text, extract the most important sentences.
[…]
Output:"
[…]
"Summarize the following input in a <TEMPLATE TYPE>style.
Style parameters: <STYLE PARAMETERS>
[…]
Output:"
Table 1: Prompts used to generate the SURe and final LaTeX document. The template is pictured in Figure 2.
3 Data
Before introducing our approach to unified generation and evaluation of templatic document views, we begin by describing the data we use in this paper.
There is no existing dataset that includes multiple views of a single document. Instead, we evaluate our unified method, described in §4, on 3 datasets: DOC2PPT, LongSumm, and Paper-Poster Fu et al. (2021); Chandrasekaran et al. (2020); Qiang et al. (2016). These datasets are chosen because they each involve generating a different view of a document. We note that our method is not specific to the scientific domain. However, the scientific domain is one of the few domains with abundantly available public data of multiple templatic views of documents. We leave exploration of other domains to future work. The three datasets and their associated generation tasks are described below.
Slide Generation.
We use use the DOC2PPT dataset introduced by Fu et al. (2021). DOC2PPT is a dataset of 5.8K scientific papers in Computer Science and their respective slide decks. We randomly sample 1K examples from this dataset for evaluation. The slides are provided as an image for each slide. We use the Azure OCR tool to extract the text from each slide [^1] https://learn.microsoft.com/en-us/azure/ai-services/computer-vision/overview-ocr.
Blog Generation.
We use the LongSumm dataset introduced by Chandrasekaran et al. (2020), which includes blog posts of scientific papers in the Computer Science domain. LongSumm was originally introduced as a shared task for the purposes of long form summarization of scientific documents. Although LongSumm includes a blind set of 22 papers, our goal is to develop both a unified method of generation and evaluation across template types. As we are unable to compute our custom evaluation metric, described in §5, we exclude the blind test data from our evaluation. Instead, we use their released train split as our evaluation data, since our method does not require any training or supervision. Of the 531 released blog posts, we were able to access 505 of their papers. The other 26 papers included broken links or were behind a paywall. Longsumm also includes a set of extractive summaries from Lev et al. (2019). We exclude these summaries from our evaluation data.
Poster Generation.
We use the Paper-Poster dataset introduced by Qiang et al. (2016). Paper-Poster is a dataset of 85 papers in Computer Science and Biology, and their respective scientific posters. We exclude 2 examples containing corrupted PDFs from our evaluation. Although Qiang et al. (2016) uses the source files to extract the text of posters for evaluation, they only release the PDF formats. To preprocess the reference posters, we found that automatic tools to extract text from documents did not handle the visual layout of posters well, so we manually extracted the text of the posters in this dataset. Note that this process was only done to obtain metrics for evaluation, and that our unsupervised generation method is capable of creating target documents without the need for reference data, or this dataset-specific manual extraction step.
For all 3 datasets, we use the Azure Document Layout tool to extract the text of the input papers. [^2] https://learn.microsoft.com/en-us/azure/ai-services/document-intelligence/concept-layout?view=doc-intel-4.0.0
4 Unified LLM-powered Generation of Templatic Views
Our goal is to develop a method of generation that is: (1) unified, (2) adaptable, and (3) requires minimal supervision. We introduce a 2 step method: generate a Structured Unified Representation (SURe) and then generate the templatic view. By first generating a SURe, we are providing the model with a structured representation of the knowledge in the input document. This representation unifies the resulting templatic views, and can be used to generate any new template styles, achieving criteria (1) and (2). Additionally, this method is a zero shot method, meaning it requires no previously annotated data or training, achieving criteria (3). This process is visualized in Figure 1.
4.1 SURe Generation
The SURe is a JSON representation of the original document, consisting of a structured layout of the most important parts of the input. We provide the document to the model along with a template of the SURe. The prompt and template used can be found in Table 1 and Figure 2, respectively. By first generating a SURe, we are providing structure to the input document, while representing only the most important information, making it easier for the LLM to parse than the full input document. While we experiment with representing the SURe as a JSON in this paper, we note that a JSON format isn’t inherent to the SURe. The SURe could be represented in any format that provides structure to the input text (e.g. XML or Markdown) and we leave exploration of different formats to future work.
4.2 Templatic View Generation
We then feed the generated SURe as input back into the LLM, asking the model to generate the final output document. We ask the LLM to represent the final output document as a LaTeX file, as LaTeX is capable of formatting multiple styles of documents, including slides, posters, and blog posts. The prompts for each step can be found in Table 1. For each templatic view, the prompt to generate the final LaTeX document takes a short description of the desired output, which we refer to as style parameters. For the tasks we evaluate in this paper, we use the following style parameters:
  • Slides: “Slides should include a title page. Following slides should contain an informative slide title and short, concise bullet points. Longer slides should be broken up into multiple slides.”
  • Posters: “Posters should include a title section at the top. Each panel should include a heading and short, concise bullet points of the most important take-aways from that section.”
  • Blogs: “Blogs should include pargraphs introducing the topic, a summary of the input document, and important takeaways. The blog should be more readable to a general audience than the input document.”
By incorporating the use of style parameters, our method is adaptable to new templatic views; the user only needs to write a short description of the template style. This means that although we evaluate our method on the scientific domain, our method is not limited and could be generalized to other domains and template styles, such as enterprise-specific document types. We leave the exploration of other domains to future work.
Figure 2: Template provided to generate the SURe.
5 Template Agnostic Evaluation
As past work has treated the generation of different templatic views as separate tasks, the evaluation of these types has also been developed independently. Unfortunately, none of the previously used metrics scale well to other document types or more generally to a unified evaluation of templatic views. ROUGE Lin (2004); Chandrasekaran et al. (2020) is only designed to evaluate text against a text reference, and does not take structure and ordering of information into account; the same is true for other fuzzy text comparison metrics Zhang* et al. (2020); Zhong et al. (2022); Yuan et al. (2021). Slide-level ROUGE Fu et al. (2021) partly solves the problem with a length penalty but still does not account for distribution of information, nor is it easily extensible to formats other than slides. Finally using a trained regressor Qiang et al. (2016) as a metric is only possible with availability of human-labeled data, which is expensive and time-consuming to curate for new domains and applications.
Thus, our goal is to develop a framework of automatic evaluation that is template agnostic. In order to generalize to multiple templates, we introduce the concept of panels. A panel is a unit of organization within a document type, for which the placement and ordering of the panel is important to the overall flow of information in the document.
For example, we consider panels to be each slide in a slide deck and each section on a poster. We consider the entirety of a blog post to be a single panel. Although we test our method on the tasks of slide, blog, and poster generation, the concept of panels is not limited to these document types. For example, each post on a social media thread could be considered a panel, or each page on a website. By generalizing to the idea of panels, we can define an evaluation metric that is template agnostic.
We aim to unify the evaluation of templatic views by adapting and integrating prior metrics into a template agnostic precision-recall framework, which we refer to as Template-Agnostic Evaluation (TAE). TAE is not a new individual metric, but rather a framework of evaluation allowing generalization to new templates. For example, although ROUGE is not designed to capture posters, TAE can be used with ROUGE to evaluate poster generation.
The general formulation of TAE is as follows:
$\begin{split}\mathrm{Precision}=Q_{P}\times O_{P}\times L\\ \mathrm{Recall}=Q_{R}\times O_{R}\times L\end{split}$ (1)
in which $Q_{P}$ is the precision measure of quality (§5.1), $O_{P}$ is the precision penalty for order (§5.2), and $L$ is the penalty for length (§5.3). Similarly $Q_{R}$ is the recall measure of quality and $O_{R}$ is the recall penalty for order. The length penalty $L$ is non-reflexive, as discussed in §5.3. We calculate separate precision and recall scores, allowing evaluators to decide which measure is most important to them, including the ability to calculate an overall F-measure score.
5.1 Quality Measure
For the TAE precision score, we calculate the average similarity between the generated panels and their most similar reference panel. More specifically, we calculate $Q_{P}$ as follows:
$Q_{P}=\frac{1}{|\tilde{S}|}\sum_{\tilde{S}}\mathrm{max}_{\mathrm{sim}}(S,\tilde{S}_{i})$ (2)
in which $S$ is the set of reference panels and $\tilde{S}$ is the set of generated panels. For the similarity metric, the user can choose a similarity metric that best matches their use case, such as ROUGE, BERT-Score, or a custom trained regressor Lin (2004); Zhang* et al. (2020). For example, a user might choose ROUGE if they want a similarity metric that focuses on exact word overlap, or they may choose BERTScore to measure broader semantic similarity.
Similar to precision, for the TAE recall score, we calculate the average similarity between the reference panels and their most similar generated panel:
$Q_{R}=\frac{1}{|S|}\sum_{S}\mathrm{max}_{\mathrm{sim}}(\tilde{S},S_{i})$ (3)
In other words, we measure how similar each reference panel is to the generated panels.
By splitting the evaluation of quality into precision and recall, we can evaluate both the content of the slides that were generated as well as the coverage of this content against some reference.
5.2 Order Penalty
Broadly, the goal of the ordering penalty is to measure the similarity of the order of information in reference and generated panels, independent of other factors. Unfortunately, because the cardinality of panels in the two outputs is not necessarily the same, a direct one-to-one mapping to compare ordering is not feasible. Instead, a panel in one set can align to multiple references in the other, or none at all – as depicted in  Figure 3. Intuitively, our solution is to virtually replicate (resp. drop) panels that have multiple (resp. zero) alignments in the reference set so that a one-to-one mapping of ordering, can in fact be computed.
Formally, assume $S$ and $\tilde{S}$ are sequences of reference and generated panels respectively. We use the maximum similarity scores calculated in §5.1 to align the panels across sets. For the precision ordering penalty, we define the following operation $\lambda_{P}(s)=\sum_{\tilde{s}}\delta_{P}(s,\tilde{s})$, where
$\delta_{P}(s,\tilde{s})=\begin{cases}1,&\text{iff }s\rightarrow\tilde{s}\\ 0,&\text{otherwise}\end{cases}$
Intuitively, this captures the cardinality of the alignment of a given panel in $S$ with panels in $\tilde{S}$. Then, using this operation we can replace every panel $s\in S$ with $\lambda(s)$ copies, leading to an identical cardinality for both $S$ and $\tilde{S}$, and subsequent one-to-one mapping between their corresponding panels.
With this one-to-one mapping in place we can then measure the misalignments between the two sets of ordered panels to assign a penalty scores. To operationalize this, we associate panels with ranks in both sets and use a rank correlation metric to compute the degree of agreement. Specifically, rank assignment is done as follows: panels in $\tilde{S}$ are simply assigned ranks in order of appearance 1 through N – we call this $\tilde{S}_{ranking}^{P}$; meanwhile panels in $S$ are assigned the identical rank to their one-to-one aligned panel in $\tilde{S}$ and $\tilde{S}_{ranking}^{P}$ – we refer to these rankings as $S_{ranking}^{P}$. The final ordering penalty is computed using Spearman’s rank correlation Szmidt and Kacprzyk (2010):
$O_{P}=\frac{\mathrm{Spearman}(S_{ranking}^{P},\tilde{S}_{ranking}^{P})+1}{2}$ (4)
where we perform a linear transformation to map the original range of the correlation coefficient [-1, 1] to the desired range [0, 1] for precision.
Similarly, for the recall ordering penalty, we perform the same operation but in the other direction, mapping the reference panels to the generated panels. This means that panels st $\delta_{P}(s,\tilde{s})=0$ will be accounted for in the recall ordering penalty. Formally, $\lambda_{R}(s)=\sum_{s}\delta_{R}(\tilde{s},s)$, where
$\delta_{R}(\tilde{s},s)=\begin{cases}1,&\text{iff }\tilde{s}\rightarrow s\\ 0,&\text{otherwise}\end{cases}$
We calculate the rankings for recall similarly, assigning the panels in $S$ a rank of $S_{ranking}^{R}=[1,...,N]$ in order of appearance, while panels in $\tilde{S}$ are assigned the identical rank to their one-to-one aligned panel in $S$ and $S_{ranking}^{R}$. We then take the Spearman’s rank correlation as follows:
$O_{R}=\frac{\mathrm{Spearman}(S_{ranking}^{R},\tilde{S}_{ranking}^{R})+1}{2}$ (5)
An example of this process can be found in Figure 3.
Figure 3: Example of the process of obtaining the rankings for the precision ordering penalty. This process is reflexive, and panels not accounted for in the precision ordering penalty are accounted for in the recall ordering penalty.
Simularity Measure
Template Model Template R-1 R-L BERTScore
slides GPT4 SURe 8.95 8.37 38.08
Base 8.89 8.29 36.24
[…]
Base 5.62 5.09 31.64
[…]
Base 27.73 25.68 82.76
[…]
Base 27.02 25.10 82.85
[…]
Base 17.61 16.35 59.82
[…]
Base 8.38 7.55 35.45
Table 2: Evaluation results using GPT3.5 (gpt35-16k) and GPT4 (gpt4-32k). For each model, we compare the performance with and without the SURe. We report the TAE F1 scores as calculated in §5, using Rouge-1 (R-1), Rouge-L (R-L), and Bert-Score as the similarity metrics.
5.3 Length Penalty
Finally, we compute a length penalty for both the recall and precision scores. Similar to Fu et al. (2021), this is done as follows:
$L=e^{\frac{-\mathrm{abs}(|S|-|\tilde{S}|)}{|S|}}$ (6)
If $L$ were reflexive, the denominator in the reverse case would be $|\tilde{S}|$, meaning that as $|\tilde{S}|\rightarrow\infty$, $L\rightarrow 1$. In other words, the length penalty could be cheated by over-generating. Therefore, unlike the quality measure and ordering penalty, we chose not to make the length penalty reflexive.
6 Results
As a baseline, we skip the SURe generation step, passing the full text of the paper as input to the generate a LaTeX document prompt, rather than the SURe. We maintain the same style parameters for both settings. Past work on the Paper-Poster dataset and the DOC2PPT dataset did not release their code or results, so we cannot directly compare to their work. LongSumm’s test set is a blind set that does not allow us to compare using our custom metric.
We use gpt4-32k and gpt35-16k for our experiments. We truncated inputs that are too long for the input window and used a temperature of 0.0, based on hyper-parameter experiments described in §B.
6.1 Results of automatic evaluation
We report the results of our automatic evaluation in Table 2, computed the TAE scores as described in §5. As we can see from the results, first generating the SURe results in performance gains for most settings. We see that largest performance gains in the gpt35-16k experiments, indicating that our method has the largest benefits for simpler models. This is likely because more simple models have more difficulty parsing long unstructured documents, and the SURe provides a structured, filtered representation of the input document. The only setting we don’t see performance gains or similar performance is in the gpt4-32k setting on the poster template. Upon closer inspection, the references in this dataset are very verbose, averaging 391 tokens total. Our method produces generally less verbose posters, averaging 265 total tokens compared to 345 tokens produced by the baseline. We hypothesize that by editing the style parameters to include information about verbosity and length, we can improve performance on posters and leave exploration of this to future work.
6.1.1 Importance of Structure
Our method provides a template JSON to the LLM to generate the SURe, before passing the resulting JSON back to the LLM to generate the final document. A natural question is, how important is the JSON format? In other words, do we see performance gains as a result of the filtering of information, or is the structure provided by the JSON important as well? In order to test this question, we conduct an additional experiment in which we take the text within the SURe, discard the JSON structure, and pass the resulting body of text to the LLM to generate the final document. We randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively. We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples. We use gpt35-16k for this experiment. The results of this experiment can be found in Table 3. As we can see from the results, passing the text of the SURe without the JSON structure decreases performance in most cases, with posters having the largest gap in performance, indicating that both the structure and content of the SURe are important for performance.
Template
SURe
Format
R-1 R-L
Bert
Score
slides JSON 7.7 7.3 35.1
Text 7.9 7.1 34.6
[…]
Text 28.3 24.4 82.5
[…]
Text 11.5 10.5 46.3
Table 3: TAE scores comparing the use of the JSON representation of the SURe or passing the text generated in the SURe, using gpt35-16k. We see that in most cases, the JSON representation performs better, indicating the importantance of structure.
Figure 4: Reasons annotators preferred each document. While annotators largely preferred documents generated with a SURe, the most common reasons for preference are better formatting and information content. We exclude the “Other” count as it was only selected once.
6.2 Human evaluation
We additionally conduct a human evaluation to test both our method and how well our evaluation framework aligns with human preference. We sample 100 documents each from Doc2PPT and LongSumm, and use the entirety of the Paper-Poster dataset. We present the annotators with 2 versions of each document, one generated with the SURe and one without. Both versions use gpt4-32k. The annotators are provided with the original paper and the intended document type (blog, slide deck, or poster), and are asked the following questions:
  1. 1.
    Which document do you prefer?
  2. 2.
    On a scale of 1-3, to what degree do you prefer your selection?
  3. 3.
    Why do you prefer your selection?
For Question 3, the annotators are given a multi-select checklist of reasons why they might prefer one document over the other. The checklist included (1) the quality of the content, (2) the formatting, (3) the document style matching the intended document types, (4) the information represented in the document, and (5) other. The “other” option included a free text box. The full annotation instructions, including a description of the reasons provided and examples, can be found in §A.
Each document was annotated by 3 different annotators. We employed 4 annotators from India, sourced via a third-party agency, to carry out the human evaluation of our task. A guideline document was provided to these evaluators, containing task-specific instructions, guidance, and several annotated examples. They were compensated at a rate of \$11.98 USD per hour for the total time that they spent working on the task, including a training round of annotation.
Which method do humans prefer?
The documents generated with a SURe were preferred by 82%, based on majority vote. 71% of the documents generated with a SURe were unanimously preferred by the annotators. We calculate a annotator agreement score of 0.51 with Krippendorff’s alpha, indicating that while this is a subjective and specialized task, even non-expert annotators agree to a moderate degree. A visualization of the reasons the annotators preferred their selection can be found in Figure 4. We can see from the results that while annotators largely preferred the documents generated with a SURe, the most common reasons for preference are better formatting and better information content. This indicates that the structure provided by the SURe makes it easier for the model to format the final document well. Additionally, the SURe only includes the most salient information from the original text, resulting in higher quality of information content. Finally, we see a fairly even distribution across the different templatic views for the reasons of preference, indicating humans prefer the documents generated with the SURe across different document types.
Metric PearsonR
TAE ROUGE-1 21.4
ROUGE-1 14.9
[…]
ROUGE-L 14.5
[…]
BERTScore 10.6*
Table 4: Correlation of evaluation metrics with human judgement. We compare each metric computed using the TAE framework versus the standard computation. *Indicates the correlation is not statistically significant ($p>0.01$)
Which metric correlates better with human judgement?
We use the human preference judgements to test if our metric, as described in §5, correlates more with human preference compared to prior evaluation metrics in the literature. For each annotation, given the degree of the preference $d$ we convert the model preference to a score $P$ as follows:
$P(d)=\begin{cases}3,&d=\textrm{strong}\\ 2,&d=\textrm{moderate}\\ 1,&d=\textrm{slight}\end{cases}$
If the annotator prefers the document generated without a SURe, we multiply $P(d)$ by $-1$. This allows us to measure if the metric captures both directionality of preference and degree.
Finally, we compute the automatic score for each metric, then calculate $S=m(\textrm{with SURe})-m(\textrm{skip SURe})$, where $m$ is the metric we are evaluating. If the annotator prefers the document generated without the SURe, we’d expect a good metric to assign a higher score to that document as well, resulting in both $S$ and $P$ being negative. Conversely, we’d expect instances where the opposite were true to result in positive scores for both $S$ and $P$. Using these intuitions we assign an affinity score of a metric with respect to human evaluation as the Pearson correlation Freedman et al. (2007) of $S$ and $P$.
Since prior metrics are not designed to account for the structure of the document, we compute them by extracting only the text of both the generated document and reference documents. The correlations with human judgement for each metric, computed standard compared to TAE variants can be found in Table 4. As we can see from the results, computing each evaluation metric using our template agnostic evaluation method correlates more highly with human judgement, except in the case of BERTScore. However, BERTScore is the only metric in which the correlation is not statistically significant, as it is likely simply a poor metric for this task. We hypothesize that BERTScore does not perform well on scientific data, since it is trained primarily on out of domain text.
7 Conclusion
In many domains, people choose to disseminate information across different modalities and formats, allowing for better communication of information to broader audiences. Past work has focused on developing methods to generate different templatic views of documents, tailoring the method to each template. In this paper, we introduced an LLM-powered method to unify the generation of templatic views of documents by first generating a Structured Unified Representation (SURe) of the input document, representing the underlying knowledge of the templatic views. We showed that our method results in improvements in performance, especially for a simpler model. We additionally introduced Template Agnostic Evaluation (TAE), a flexible precision-recall framework for automatic evaluation that is template agnostic, allowing for easy integration of existing evaluation metrics into a unified system. The framework is capable of jointly accounting for and measuring multiple facets of the output document including the quality of its content, its underlying structure and ordering, as well as its length. Finally, we conducted a human evaluation and showed that annotators prefer the documents generated using the SURe 82% of the time. We additionally showed that our framework of evaluation correlates more highly with human preference than standard evaluation metrics.
The use of templatic views of documents exists in many domains (e.g. creating a cover letter and personal website based on a resume), and in future work we will consider other domains of documents, as well as more complex scenarios, such as many-to-many document generation. Our work is a step towards a larger goal of generalized document generation.
Limitations
Although our methods are not domain specific, we only evaluated our methods in the scientific domain, due to the availability of public data. Further research is required to evaluate the performance of our methods in other domains. Additionally, our method is limited to textual content. Future work should explore incorporating multi-modal models and content generation. Finally, we evaluated our methods on closed source models, which may affect the reproducibility of our results.
Appendix A Annotation Instructions
A.1 Questions
Question 1 – Which document do you prefer? In this question, you are asked to choose which document version you prefer. Some examples of qualities you may use to decide your preference include:
  • The quality of the content – The text is grammatical and understandable. E.g. Document A contains major grammatical errors while Document B only contains minor errors.
  • The formatting – The formatting is reasonable and matches the formatting of the intended document type. E.g. A poster contains panels and each panel contains a header and body text.
  • The style – The document matches the style of the intended document type. E.g. Shorter, bulleted sentences in a slide deck.
  • Information represented in the document – The document contains sufficient information to represent the input document. E.g. A blog post represents the most important sections from the input document.
The above criteria are non-exhaustive. Not all criteria must be met, and you may use other relevant criteria to make your decision. You are not rating the document for factual correctness, and only need to refer to the corresponding scientific article if it will aid in making your preference. You can answer this question with either Document A or Document B.
Question 2 – On a scale of 1-3, to what degree do you prefer your selection?
In this question you will rate the degree to which you prefer your selection, on the following scale:
  1. 1.
    Small preference – The documents are similar in quality and only contain minor differences that affect my preference.
  2. 2.
    Moderate preference – I clearly prefer one document but the differences are not major.
  3. 3.
    Strong preference – I have a strong preference for one document and the differences between the documents are major.
Question 3 – Why do you prefer your selection? (You may select more than one property)
  • $\square$
    Formatting
  • $\square$
    Information
  • $\square$
    Quality
  • $\square$
    Style
  • $\square$
    Other (free text)
A.2 Edge cases
For most edge cases, it is up to your discretion on how to best handle the case. However, below are a few examples of how you could consider certain edge cases:
Example 1: Slides 1-5 of Document A are higher quality but slides 6-10 of Document B are higher quality. You could reason that the first slides represent the most important information, and choose Document A. However, since Document B contained higher quality slides for another portion of the document, you could rate your degree of preference as “Small preference.”
Example 2: Document A more closely matches the style of the intended document type, but Document B contains more relevant information to the source document. You could consider if Document A contains sufficient information to represent the input document, such as representing the most important sections. If yes, then you could prefer Document A. If not, then you could reason that information content is more important than style, and prefer Document B.
Example 3: Document A contains more relevant information than Document B, but also contains major formatting errors, such as text being cut off from the document.
You could reason that although Document A contains more relevant information, the major formatting errors are significant enough to prefer Document B.
Example 4: Neither document matches the style or formatting of the intended document type. Since neither document matches the style or formatting of the intended document type, you could consider other criteria, such as quality of content or the information represented.
Template Temp R-1 R-L BertScore
slides 0 7.7 7.3 35.1
0.25 7.8 7.2 35.4
0.5 7.6 7.0 36.4
0.75 7.5 7.0 35.5
1 8.0 7.4 35.8
[…]
0.25 27.2 25.5 82.6
0.5 28.0 26.2 82.8
0.75 27.2 25.4 82.6
1 26.6 24.8 82.7
[…]
0.25 14.2 13.0 53.1
0.5 13.6 12.5 52.2
0.75 13.2 12.0 50.8
1 12.6 11.6 50.3
Table 5: Results of the temperature hyperparameter experiments. There seems to be little consistency in performance as related to temperature across different templates.
Appendix B Temperature Experiments
We experiment with the temperature of the generations to see how temperature affects performance. Similar to the textual representation experiments, described in §6.1.1, we randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively. We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples. We use gpt35-16k and experiment with the temperatures $[0.0,0.25,0.5,0.75,1.0]$. The results of this experiment can be found in Table 5. As we can see from the results, there seems to be little consistency across the different types in which temperature performs the best.