### v2

# Knowledge-Centric Templatic Views of Documents

Abstract
Authors seeking to communicate with broader audiences often share their ideas in various document formats, such as slide decks, newsletters, reports, and posters. Prior work on document generation has generally tackled the creation of each separate format to be a different task, leading to fragmented learning processes, redundancy in models and methods, and disjointed evaluation. We consider each of these documents as templatic views of the same underlying knowledge/content, and we aim to unify the generation and evaluation of these templatic views. We begin by showing that current LLMs are capable of generating various document formats with little to no supervision. Further, a simple augmentation involving a structured intermediate representation can improve performance, especially for smaller models. We then introduce a novel unified evaluation framework that can be adapted to measuring the quality of document generators for heterogeneous downstream applications. This evaluation is adaptable to a range of user defined criteria and application scenarios, obviating the need for task specific evaluation metrics. Finally, we conduct a human evaluation, which shows that people prefer 82% of the documents generated with our method, while correlating more highly with our unified evaluation framework than prior metrics in the literature.
Figure 1: Visualization of our method to unify the generation and evaluation of templatic views of documents. Given an input document, we prompt the LLM to generate an intermediate representation. We can use the representation to prompt the model to generate a templatic view of the input document. We then evaluate the generations using our unified evaluation framework. The LLM represented in the figure is the same model.
1 Introduction
Sharing information is vital for communication and discourse across domains, as it allows for knowledge to be disseminated to a wider audience. This is often done by users through documents in multiple formats that nevertheless share some underlying knowledge. A product manager may need to create a requirements spec, a product pitch deck, and an announcement newsletter for the same project. Likewise, a person on the job market may create a resume, a cover letter, and a personal website. We consider these documents to be templatic views of the same underlying knowledge.
This is equally true for the scientific domain, in which researchers create documents in multiple formats to effectively communicate and showcase their work, – such as through academic papers, conference talks, social media posts, poster presentations, and non-technical blog posts. Sharing knowledge in multiple formats broadens the audience and can help bridge the information gap between domain experts, researchers in adjacent fields, and even the general public, leading to greater understanding, collaborations and accelerated progress Bornmann and Mutz (2014).
Past work on document generation has focused on developing generation and evaluation methods specific to a single document type Fu et al. (2021); Qiang et al. (2016); Chandrasekaran et al. (2020). Narrow, custom methods tailored to individual document types are, nevertheless, time consuming to engineer and manage over the long term. For example, in an enterprise setting, it’s common to have dozens of occupation- and task-specific documents, each with their own template.Additionally, specific trained methods require data that may be expensive to acquire, or even be unavailable entirely. Meanwhile, LLMs have recently shown great success in long document generation Radford et al. (2019); Brown et al. (2020), indicating that this fragmentation of methods may no longer be necessary. Thus, our goal is to unify methods for both generating and evaluating templatic views of documents, allowing system designers and engineers to manage and adapt to a range of document types and domains easily and efficiently.
We begin by showing that LLMs are capable of diverse, structured document generation, requiring very little instructional guidance to do so effectively. Additionally, a few minor augmentations to the prompt – such as a structured, intermediate representation, and simple stylistic descriptions – can further improve downstream performance, especially for smaller, less resource intensive models. These findings have important implications on the deployment and scaling of unified, real-world AI-assisted document authoring systems.
In similar vein, we then introduce Template Adaptable Evaluation (TAE), departing from prior work’s task specific evaluation methods Zhang* et al. (2020); Qiang et al. (2016); Wang et al. (2015). TAE is a unified precision-recall style framework for automatic evaluation that is highly customizable, allowing users to easily integrate existing text-based metrics from the literature into its formulation and tailor it to their specific use case.Additionally, this framework allows developers to compare performance across document types, without needing to develop an evaluation metric for each individual template.
We evaluate our unified approach for templatic view generation and evaluation on 3 types of documents: slides, posters, and blog posts Fu et al. (2021); Qiang et al. (2016); Chandrasekaran et al. (2020). Our experiments demonstrate that using a structured intermediate representation leads to improvements in performance across tasks, with greater gains for smaller language language models. In our human evaluation to validate both our unified document generation method and evaluation metric, we show that annotators prefer the output yielded by the structure-aware generation process 82% of the time and that our evaluation metric correlates more highly with human preference than other popular metrics. We release our code [^1] https://github.com/microsoft/knowledge-centric-templatic-views to support future research.
2 Related Work
There are several areas of related research in NLP that are relevant to the problems of document transformation and evaluation.
Document summarization has been explored in a number of domains, including news See et al. (2017), literature Sciré et al. (2023), law Deroy et al. (2023), and dialogue Chen et al. (2021). In the scientific domain, summarization of scientific papers has taken the form of long form summaries Chandrasekaran et al. (2020), abstract generation Cohan and Goharian (2015), conference talks Lev et al. (2019), and query based summaries Fok et al. (2023). These summaries can be either extractive Sefid and Giles (2022) or abstractive Chandrasekaran et al. (2020).
Although the tasks of slide and poster generation have generally been considered separate from scientific summarization, they are related in that both tasks require taking an input article, then organizing and abstracting the information to generate a new document. Past work has developed methods for slide generation from papers Hu and Wan (2015); Li et al. (2021); Hu and Wan (2015); Fu et al. (2021), from code Wang et al. (2023a), or based on a query Sun et al. (2021). Poster generation has been explored in the form of content extraction for posters Xu and Wan (2021), interactive generation Wang et al. (2015), or full content generation using graphical models Qiang et al. (2016). To the best of our knowledge, our work is the first to create a unified method capable of generating a diverse range of templatic views of a source document.
Large Language Models (LLMs), which are central to our approach, have shown impressive capabilities in a variety of tasks Radford et al. (2019); Brown et al. (2020). Based on the transformer architecture Vaswani et al. (2017), LLMs have shown emergent abilities in tasks such as arithmetic and question answering Wei et al. (2022a). Similar to chain of thought prompting Wei et al. (2022b) and content planning prompting Wang et al. (2023b), we show that by generating an intermediate representation of an input document can improve performance over simply prompting the model to generate the final document from the original input.
As past work has tackled generation of templatic views as separate tasks, methods for automatic evaluation of different document types is fragmented. LongSumm, the shared task introduced by Chandrasekaran et al. (2020), uses ROUGE to evaluate model performance Lin (2004). Fu et al. (2021) introduced Slide Level ROUGE to evaluate slide generation, a variant that contains a penalty for the number of slides. Qiang et al. (2016) used a trained regressor. For summarization, many automatic evaluation metrics have been introduced such as BERTScore Zhang* et al. (2020), UniEval Zhong et al. (2022), BARTScore Yuan et al. (2021), BLANC Vasilyev et al. (2020), and MoverScore Zhao et al. (2019). However, these metrics are intended for a simple input document-summary setup, and do not take into account factors that affect the quality of other types of documents (e.g. structure). Our work is the first to introduce template adaptable evaluation, allowing uniform comparison of performance across template types.
3 Data
We begin by describing the data used in this paper. There is no existing dataset that includes multiple views of a single document. Instead, we evaluate our unified method, described in §4, on 3 existing datasets: DOC2PPT, LongSumm, and Paper-Poster Fu et al. (2021); Chandrasekaran et al. (2020); Qiang et al. (2016). These datasets are chosen because they each involve generating a different view of a document. Although our method is not specific to the scientific domain, it is one of the few domains with abundantly available public data of multiple templatic views  [^2] We acknowledge that scientific writing does have structural regularities that may influence unified document generation. Due to the lack of other available datasets we leave exploration of other domains to future work.. The three datasets and their associated generation tasks are described below.
Slide Generation.
We use use the DOC2PPT dataset Fu et al. (2021), which contains 5.8K scientific papers in Computer Science and their respective slide decks. As Fu et al. (2021) do not release data splits or code, we randomly sample 1K examples from this dataset for evaluation. The slides are provided as an image for each slide. We use the Azure OCR tool to extract the text from each slide [^3] https://learn.microsoft.com/en-us/azure/ai-services/computer-vision/overview-ocr.
Blog Generation.
We use the LongSumm dataset Chandrasekaran et al. (2020), which includes blog posts of scientific papers in the Computer Science domain. Since our approach requires no training or supervision, we use the entire training split from Longsumm as our evaluation set. Of the 531 publicly released blog posts in this set, we could only access 505, with the other 26 including broken links or being behind a paywall.
Notably, while Longsumm includes a blind test set of 22 papers, this test set only consists of inputs without their reference outputs, thus making it impossible to compute our custom evaluation metric (see §5). In the interest of completeness and comparison to prior work, we do, however submit runs from our systems to the leader board and report the results of this blind test set in Appendix D.
Poster Generation.
We use the Paper-Poster dataset Qiang et al. (2016), which consists of a dataset of 85 papers in Computer Science and Biology, and their respective scientific posters; two examples containing corrupted PDFs are excluded. Although Qiang et al. (2016) release data splits, they do not release code or results for comparison. Given the small size of the dataset, we use it in its entirety for more robust results. While the authors uses the source files to extract the text of posters for evaluation, they only release the PDF formats. To preprocess the reference posters, we found that automatic tools to extract text from documents did not handle the visual layout of posters well, so we manually extracted the text of the posters in this dataset. Note that this process was only done to obtain evaluation scores, and that our unsupervised generation method is capable of creating target documents without the need for reference data.
For all 3 datasets, we use the Azure Document Layout tool to extract the text of the input papers. [^4] https://learn.microsoft.com/en-us/azure/ai-services/document-intelligence/concept-layout
4 Unified LLM-powered Generation of Templatic Views
The most straightforward way to transform documents between templatic views using LLMs, is to simply prompt the system to generate the target view given the input. However, similar to chain of thought prompting Wei et al. (2022b), we hypothesize that first generating a structured, intermediate representation of an input document and then reasoning over that representation will result in better generations than directly prompting the model. Our goal is to evaluate the capabilites of LLMs to generate long, structured documents, and experiment with how structured prompting can improve performance. We experiment with a simple general two-step process: first generate an intermediate representation, then generate the templatic view. These steps are described in greater detail below, and the process is visualized in Figure 1.
Intermediate Representation Generation.
In this work, we set the intermediate representation to be a JSON consisting of a structured layout of the most important parts of the input. We provide the input document to the model along with a template of the representation and prompt it to extract the most important information from the input document, and format it in the given JSON structure. The exact prompts and JSON structure can be found in Appendix §A. While our experiments use a JSON intermediate representation, note that other formats that provide structure to the input text could be employed (e.g. XML or Markdown). Rather than trying to optimize for the best representation format, our goal is to show that this chain of extraction approach along with structured augmentation to prompts can aid the quality of generations from LLMs. We leave exploration of different formats and other prompt optimization to future work.
Templatic View Generation.
We then feed the generated representation as input back into the LLM, prompting the model to generate the final output document, represented as a LaTeX document. For each templatic view, the prompt to generate the final LaTeX document takes a short description of the desired output, which we refer to as a style parameter. For example, the style parameter for slide generation is as follows: “Slides should include a title page. Following slides should contain an informative slide title and short, concise bullet points. Longer slides should be broken up into multiple slides.” The use of style parameters makes our method adaptable to new templatic views; the user only needs to write a short description of the template style. Both the generation of the intermediate representations and the final documents require little to no prompt engineering. The prompts and style parameters can be found in Appendix §A.
5 Template Adaptable Evaluation
Prior work on document generation has treated the evaluation of different templatic views as separate tasks. Thus, our goal is to develop a framework of automatic evaluation that is template adaptable. This not only allows us to compare performance across diverse datasets, it also removes the requirement of designing and maintaining individual metrics for each template. In order to generalize to multiple templates, we introduce the concept of panels. A panel is a unit of organization within a document type, for which the placement and ordering of the panel is important to the overall flow of information in the document.
For example, we consider panels to be each slide in a slide deck and each section on a poster. We consider the entirety of a blog post to be a single panel. Although we test our method on the tasks of slide, blog, and poster generation, the concept of panels is not limited to these document types. For example, each post on a social media thread could be considered a panel, or each page on a website.
We aim to unify the evaluation of templatic views by integrating prior metrics into a template adaptable precision-recall framework, which we refer to as Template-Adaptable Evaluation (TAE). TAE is not a new individual metric, but rather an evaluation framework that allows generalization to new templates. For example, TAE can even be used with ROUGE to evaluate poster generation.
The general TAE formulation is as follows:
$\begin{split}\mathrm{Precision}=Q_{P}\times O_{P}\times L\\ \mathrm{Recall}=Q_{R}\times O_{R}\times L\end{split}$ (1)
in which $Q_{P}$ is the precision measure of quality (§5.1), $O_{P}$ is the precision penalty for order (§5.2), and $L$ is the non-reflexive penalty for length (§5.3). Similarly, $Q_{R}$ is the recall quality measure and $O_{R}$ is the recall penalty for order. The precision-recall formulation allows evaluators to decide which measure is most important to them, or calculate an overall F-measure score.
5.1 Quality Measure
For the TAE precision score, we calculate the average similarity between the generated panels and their most similar reference panel as follows:
$Q_{P}=\frac{1}{|\tilde{S}|}\sum_{\tilde{S}}\mathrm{max}_{\mathrm{sim}}(S,\tilde{S}_{i})$ (2)
in which $S$ is the set of reference panels and $\tilde{S}$ is the set of generated panels. For the similarity metric, the user can choose a metric that best matches their use case, such as ROUGE, BERT-Score, or a custom trained regressor Lin (2004); Zhang* et al. (2020). For example, a user might choose ROUGE if they want a similarity metric that focuses on exact word overlap, or BERTScore to measure broader semantic similarity.
Similar to precision, the TAE recall score is calculated as the average similarity between the reference panels and their most similar generated panel:
$Q_{R}=\frac{1}{|S|}\sum_{S}\mathrm{max}_{\mathrm{sim}}(\tilde{S},S_{i})$ (3)
By splitting the evaluation of quality into precision and recall, we can evaluate both the content of the slides that were generated as well as the coverage of this content against some reference.
5.2 Order Penalty
Broadly, the goal of the ordering penalty is to measure the similarity of the order of information in reference and generated panels, independent of other factors. Unfortunately, because the cardinality of panels in the two outputs is not necessarily the same, a direct one-to-one mapping to compare ordering is not feasible. Instead, a panel in one set can align to multiple references in the other, or none at all – as depicted in  Figure 2. Intuitively, our solution is to virtually replicate (resp. drop) panels that have multiple (resp. zero) alignments in the reference set so that a one-to-one mapping of ordering, can in fact be computed.
Formally, assume $S$ and $\tilde{S}$ are sequences of reference and generated panels respectively. We use the maximum similarity scores calculated in §5.1 to align the panels across sets.
For the precision ordering penalty, we define the following operation $\lambda_{P}(s)=\sum_{\tilde{s}}\delta_{P}(s,\tilde{s})$, where
$\delta_{P}(s,\tilde{s})=\begin{cases}1,&\text{iff }s\rightarrow\tilde{s}\\ 0,&\text{otherwise}\end{cases}$
Intuitively, this captures the cardinality of the alignment of a panel in $S$ with panels in $\tilde{S}$. Then, using this operation we can replace every $s\in S$ with $\lambda_{P}(s)$ copies, leading to an identical cardinality for both $S$ and $\tilde{S}$, and subsequent one-to-one mapping between their corresponding panels.
Figure 2: Example of the process of obtaining the rankings for the precision ordering penalty. We first use the similarity measure to map each generated panel to its most similar reference document. This mapping is used to calculate the precision quality score $Q_{P}$. We then use the mappings to create a one-to-one alignment from the generated to the reference panels, which we use to calculate the precision ordering penalty ($O_{P}$). By creating a one-to-one alignment, we are able to represent inversions in the ordering. This process is reflexive, and panels not accounted for in the precision ordering penalty are accounted for in the recall ordering penalty.
Then, to operationalize a penalty score for the two sets of ordered panels we associate them with ranks in both sets and use a rank correlation metric to compute the degree of agreement. Specifically, rank assignment is done as follows: panels in $\tilde{S}$ are simply assigned ranks in order of appearance 1 through N – we call this $\tilde{S}_{ranking}^{P}$; meanwhile panels in $S$ are assigned the identical rank to their one-to-one aligned panel in $\tilde{S}$ and $\tilde{S}_{ranking}^{P}$ – we refer to these rankings as $S_{ranking}^{P}$. An example of this process can be found in Figure 2. The final ordering penalty is computed using Spearman’s rank correlation Szmidt and Kacprzyk (2010):
$O_{P}=\frac{\mathrm{Spearman}(S_{ranking}^{P},\tilde{S}_{ranking}^{P})+1}{2}$ (4)
where we perform a linear transformation to map the original range of the correlation coefficient [-1, 1] to the desired range [0, 1].
Similarly, for the recall ordering penalty, we map the reference panels to the generated panels, calculated as $\lambda_{R}(s)=\sum_{s}\delta_{R}(\tilde{s},s)$. $O_{R}$ is calculated similar to $O_{P}$, using the recall rankings.
5.3 Length Penalty
Finally, we compute a length penalty for both the recall and precision scores. Similar to Fu et al. (2021), this is done as follows:
$L=e^{\frac{-\mathrm{abs}(|S|-|\tilde{S}|)}{|S|}}$ (5)
We chose to keep $L$ non-reflexive, because in the reverse case – as $|\tilde{S}|\rightarrow\infty$, $L\rightarrow 1$ – the metric could be cheated by over-generating.
6 Results
As mentioned in §3, past work on Doc2PPT and Paper-Posters do not release code, making it difficult to do a direct comparison. They also do not report any baselines to compare against. Meanwhile, Longsumm’s blind test does not allow us to compute our custom metric, although we do report the leaderboard results in Appendix D. Notably, with almost no prompt engineering our LLM-based system places second on this leaderboard. We argue that for the investigation in this paper, direct comparison to prior non-LLM baselines is not only unfair to those approaches, but not particularly insightful. Therefore, similar to Wei et al. (2022b), we focus on variants of our LLM-based method and treat them as baselines. Example outputs of each template type can be found in Appendix E.
We conduct experiments with the following settings: (1) No Representation – this is the default setting of going directly from the source document to the target document. We skip the intermediate generation step, passing the full paper as input. We experiment both with and without the style parameters. (2) Own Representation – we do not pass a JSON structure to the intermediate generation step, and allow the model to choose its own structure. (3) Text Representation – we extract the text from the intermediate representation, discarding the JSON structure. (4) JSON Representation – this is the full JSON structure for the intermediate generation step. We experiment both with and without the style parameters.
We use gpt35-16k in our main set of experiments. We truncate text that is too long for the input window and use a temperature of 0.0 as standard. [^5] A detailed evaluation of the temperature hyper-parameter is included in Appendix §C
Similarity Measure
Rep. Style R-L M B BERTS
Slides None $\times$ 5.0 6.4 0.3 31.6
None $\checkmark$ 5.1 6.0 0.4 31.7
Own $\checkmark$ 6.5 7.1 1.2 36.1
Text $\checkmark$ 7.3 8.0 1.4 36.4
JSON $\times$ 4.2 6.0 0.3 31.4
JSON $\checkmark$ 7.4 8.4 1.5 36.9
[…]
None $\checkmark$ 25.1 17.7 2.3 82.8
Own $\checkmark$ 23.9 19.2 2.3 82.2
Text $\checkmark$ 25.4 19.3 2.5 82.5
JSON $\times$ 28.3 25.3 5.0 82.3
JSON $\checkmark$ 25.4 19.6 2.8 82.4
[…]
None $\checkmark$ 10.1 11.6 1.9 39.5
Own $\checkmark$ 12.8 12.6 2.9 52.8
Text $\checkmark$ 11.3 11.7 2.1 45.9
JSON $\times$ 14.2 16.8 4.0 52.8
JSON $\checkmark$ 15.5 14.5 15.3 53.3
Table 1: Evaluation results using GPT3.5 (gpt35-16k). For each template, we experiment with different representations (Rep) and whether or not we include the style parameters (Style). We report the TAE F1 scores as calculated in §5, using ROUGE-L (R-L), METEOR (M), BLEU (B), and BERTScore (BERTS) as the similarity metrics.
6.1 Results of automatic evaluation
In Table 1, we report the TAE F1 scores as described in §5, using ROUGE-L Lin (2004), METEOR Banerjee and Lavie (2005), BLEU Papineni et al. (2002), and BERTScore Zhang* et al. (2020) for the similarity measure. As seen in the results, by most measures, generating a JSON intermediate representation yields the best performance.
We see that using the text representation generally degrades the performance over providing the structured JSON representation, indicating that structure is important for downstream performance in addition to abstractive filtering of information. Additionally, the text representation performs better than skipping the intermediate step altogether for both the poster and slide generation task, but not the blog generation task. This is likely because posters and slides have more inherent structure than blog posts, which can be relatively free-form.
Finally, we see that allowing the model to choose its own representation format degrades performance over providing our pre-defined JSON structure. However, we see that in most cases, providing a representation generated without a JSON structure still performs better than skipping the intermediate generation step altogether (while maintaining the same style parameter setting). This indicates that even without a pre-defined structure, the intermediate step is still valuable for performance.
Similarity Measure
Model Rep. R-L M B BERTS
Slides MS $\times$ 0.6 0.4 0.0 28.6
$\checkmark$ 4.4 4.6 0.4 30.5
[…]
$\checkmark$ 6.7 7.9 1.0 34.0
[…]
$\checkmark$ 8.4 9.1 2.0 38.1
[…]
$\checkmark$ 21.7 16.2 1.7 81.3
[…]
$\checkmark$ 25.6 20.5 3.2 82.5
[…]
$\checkmark$ 25.8 20.2 3.1 82.7
[…]
$\checkmark$ 6.0 6.5 1.2 38.1
[…]
$\checkmark$ 10.4 11.0 1.5 50.7
[…]
$\checkmark$ 14.6 15.3 3.7 57.2
Table 2: TAE F1 scores using Mistral-7b (MS), Mixtral (MX), and GPT4. We use ROUGE-L (R-L), METEOR (M), BLEU (B) and BERTScore (BERTS) as our similarity measures. For each template, we compare a JSON representation versus skipping the intermediate generation step (Rep), maintaining the same style parameters in both settings.
Experiments with additional models.
[…]
Figure 3: Reasons annotators preferred each document. While annotators largely preferred documents generated with an intermediate representation, the most common reasons for preference are better formatting and information content. We exclude the “Other” count as it was only selected once.
6.2 Human evaluation
After showing that LLMs benefit from intermediate structured representations in document transformations, we investigate whether our proposed evaluation framework aligns better with human judgment than previously proposed metrics. We sample 100 documents each from DOC2PPT and LongSumm, and use the entirety of the Paper-Poster dataset in this study. We present annotators with 2 versions of each document, one generated with the intermediate representation and one without. Both versions use gpt4-32k, as the best performing model.
The annotators are provided with the original paper and the intended document type (blog, slide deck, or poster), and are asked the following questions: (1) Which document do you prefer? (2) On a scale of 1-3, to what degree do you prefer your selection? (3) Why do you prefer your selection? For question 3, annotators are also provided with a multi-select checklist of reasons for their preference: (1) quality of the content, (2) formatting, (3) document style matching the intended document types, (4) information represented in the document, and (5) other (along with a free text box). The full instructions, including the reasons provided and examples, can be found in Appendix §B.
[…]
Which method do humans prefer?
The documents generated with an intermediate representation were preferred by 82%, based on majority vote (71% unanimously). The annotator agreement score was 0.51 with Krippendorff’s alpha, indicating that while this is a subjective and specialized task, even non-expert annotators agree to a moderate degree. A visualization of the reasons the annotators preferred their selection can be found in Figure 3. It can be seen that while annotators largely preferred the documents generated with an intermediate representation, the most common reasons for preference are better formatting and better information content. This indicates that the structure provided by the intermediate representation makes it easier for the model to format the final document well. Additionally, the intermediate representation only includes the most salient information from the original text, resulting in higher quality of information content. Finally, we see a fairly even distribution across different templatic views for the reasons of preference, indicating humans prefer the documents generated with the intermediate representation across different document types.
Which metric correlates better with humans?
We test whether our metric, as described in §5, correlates better with human preference compared to prior evaluation metrics in the literature. For each annotation, given the degree of the preference $d$ (Appendix §B Q. 2) we convert value to a score $P(d)\rightarrow[1,2,3]$ if $d$ is slight, moderate, or strong, respectively. If the annotator prefers the document generated without an intermediate representation, we take $-P(d)$ instead. This allows us to measure if the metric captures directionality of preference along with degree. in parallel, we compute the automatic score $m$ for each metric, then calculate $S=m(\textrm{with rep})-m(\textrm{skip rep})$ where $m$ is the metric we are evaluating (e.g ROUGE). If a human annotator prefers a document generated without the intermediate step, we’d expect a good metric to assign a higher score to that document as well, resulting in both $S$ and $P(d)$ being negative (and positive in the opposite case). Using this intuition we assign an affinity score of a metric with respect to human evaluation as the Pearson correlation Freedman et al. (2007) of $S$ and $P(d)$.
Since prior metrics are not designed to account for the structure of documents, we compute them by extracting only the text of both the generated and reference documents. The correlations with human judgement for each metric to its respective TAE variants can be found in Table 3. As we can see from the results, evaluations using our template adaptable framework correlate more highly with human judgement, except in the case of BERTScore. In the latter case the results are not statistically significant, and we hypothesize that the open-domain nature of BERT embeddings are poorly suited to represent the semantic similarity of scientific text.
Metric PearsonR
ROUGE-L 14.5
TAE ROUGE-L 19.7
[…]
TAE METEOR 25.2
[…]
TAE BLEU 13.8
[…]
TAE BERTScore 5.4*
Table 3: Correlation of evaluation metrics with human judgement. We compare each metric computed using the TAE framework versus the standard computation. *Indicates the correlation is not statistically significant ($p>0.01$).
7 Conclusion
In many domains, people choose to disseminate information across different modalities and formats for better communication to broader audiences. We proposed a unified view of document transformation and evaluation. We showed that LLMs are capable of templatic document generation with minimal supervision, and that a structured, intermediate representation can improve performance, particularly for smaller models. We also introduced a flexible precision-recall framework for automatic evaluation that easily integrates existing evaluation metrics into a unified system and allows for comparison across diverse datasets without additional task specific metric design. Finally, we conducted a human evaluation and showed that annotators prefer the documents generated using the intermediate representation 82% of the time and that our evaluation framework correlates better with human preference than standard evaluation metrics.
8 Limitations
Although our methods are not domain specific, we only evaluated them in the scientific domain, due to the availability of public data. Additionally, our framework is limited to textual content. In future work we plan to explore the application of our unified framework for generation and evaluation on document views in other domains, as well as incorporating multi-modal models and content generation. Finally, it is possible that some of our test data has leaked into the training data of the models with which we experimented. This limitation is not unique to our work and exists for our baselines in addition to our methods.
9 Ethics
The potential risks of our work are similar to those of other work in downstream applications of LLMs. LLM generated documents can potential generate copy-righted material Carlini et al. (2020), personally-identifiable information Lukas et al. (2023), or factually incorrect text Manakul et al. (2023). The use of LLMs to generate documents may violate some academic dishonesty policies Zdravkova et al. (2023). Our system is intended to be used in collaboration with human writers. Users should edit the generations, checking for factual inconsistencies and other potential errors. Our work is intended to save users time that might be spent repeating information across multiple documents, so they can focus on content creation. Therefore, we believe the benefits of our work outweigh the potential risks.
Appendix A Prompt details
The prompts and intermediate representation template used can be found in Table 4 and Figure 4, respectively. We note that the specific structure provided to the prompt is not inherent to our method, and a different structure could be provided depending on the input document and domain. For the tasks we evaluate in this paper, we use the following style parameters:
  • Slides: “Slides should include a title page. Following slides should contain an informative slide title and short, concise bullet points. Longer slides should be broken up into multiple slides.”
  • Posters: “Posters should include a title section at the top. Each panel should include a heading and short, concise bullet points of the most important take-aways from that section.”
  • Blogs: “Blogs should include paragraphs introducing the topic, a summary of the input document, and important takeaways. The blog should be more readable to a general audience than the input document.”
Figure 4: Template of the intermediate representation provided to the prompts in Table 4.
Appendix B Annotation Instructions
B.1 Questions
Question 1 – Which document do you prefer? In this question, you are asked to choose which document version you prefer. Some examples of qualities you may use to decide your preference include:
  • The quality of the content – The text is grammatical and understandable. E.g. Document A contains major grammatical errors while Document B only contains minor errors.
  • The formatting – The formatting is reasonable and matches the formatting of the intended document type. E.g. A poster contains panels and each panel contains a header and body text.
  • The style – The document matches the style of the intended document type. E.g. Shorter, bulleted sentences in a slide deck.
  • Information represented in the document – The document contains sufficient information to represent the input document. E.g. A blog post represents the most important sections from the input document.
The above criteria are non-exhaustive. Not all criteria must be met, and you may use other relevant criteria to make your decision. You are not rating the document for factual correctness, [^6] The annotators are non-experts and do not have the background to determine factual correctness of scientific information. Instead, they are encouraged to use the original paper to understand if the information presented in the documents represent the information in the paper, to the best of their understanding. and only need to refer to the corresponding scientific article if it will aid in making your preference. You can answer this question with either Document A or Document B.
Question 2 – On a scale of 1-3, to what degree do you prefer your selection?
In this question you will rate the degree to which you prefer your selection, on the following scale:
  1. 1.
    Small preference – The documents are similar in quality and only contain minor differences that affect my preference.
  2. 2.
    Moderate preference – I clearly prefer one document but the differences are not major.
  3. 3.
    Strong preference – I have a strong preference for one document and the differences between the documents are major.
Question 3 – Why do you prefer your selection? (You may select more than one property)
  • $\square$
    Formatting
  • $\square$
    Information
  • $\square$
    Quality
  • $\square$
    Style
  • $\square$
    Other (free text)
B.2 Edge cases
For most edge cases, it is up to your discretion on how to best handle the case. However, below are a few examples of how you could consider certain edge cases:
Example 1: Slides 1-5 of Document A are higher quality but slides 6-10 of Document B are higher quality. You could reason that the first slides represent the most important information, and choose Document A. However, since Document B contained higher quality slides for another portion of the document, you could rate your degree of preference as “Small preference.”
Example 2: Document A more closely matches the style of the intended document type, but Document B contains more relevant information to the source document. You could consider if Document A contains sufficient information to represent the input document, such as representing the most important sections. If yes, then you could prefer Document A. If not, then you could reason that information content is more important than style, and prefer Document B.
Example 3: Document A contains more relevant information than Document B, but also contains major formatting errors, such as text being cut off from the document.
You could reason that although Document A contains more relevant information, the major formatting errors are significant enough to prefer Document B.
Example 4: Neither document matches the style or formatting of the intended document type. Since neither document matches the style or formatting of the intended document type, you could consider other criteria, such as quality of content or the information represented.
Simularity Measure
Temp R-L M B BERTS
Slides 0.0 7.3 8.2 1.6 35.1
0.25 7.2 8.3 1.5 35.4
0.5 7.0 8.3 1.5 36.4
0.75 7.0 8.0 1.2 35.5
1.0 7.4 8.2 1.4 35.8
[…]
0.25 25.5 19.8 2.7 82.6
0.5 26.2 20.9 3.2 82.8
0.75 25.4 19.9 2.7 82.6
1.0 24.8 19.9 2.6 82.7
[…]
0.25 13.0 14.8 3.4 53.1
0.5 12.5 14.0 2.7 52.2
0.75 12.0 13.9 3.0 50.8
1.0 11.6 11.9 2.4 50.3
Table 5: Results of the temperature hyperparameter experiments. We use ROUGE-L (R-L), METEOR (M), BLEU (B) and BERTScore (BERTS) as our similarity measures.
Appendix C Temperature Experiments
We experiment with the temperature of the generations to see how temperature affects performance. We randomly sample 100 documents each from LongSumm and Doc2PPT for the blog and poster generation tasks, respectively. We use the entirety of the Paper-Poster dataset, since it contains less than 100 examples. We use gpt35-16k and experiment with the temperatures $[0.0,0.25,0.5,0.75,1.0]$. The results of this experiment can be found in Table 5. As we can see from the results, there seems to be little consistency across the different types in which temperature performs the best.
Appendix D Longsumm Blind Test Set Results
We submit the final documents from GPT4, the best performing model overall, to the Longsumm blind test set evaluation. We compare the documents generated with and without the intermediate step. We see that without the intermediate representation we get a Rouge-1 score of 46.8 while the results generated without the intermediate representation received a Rouge-1 score of 46.4. We note that this blind test set of 22 papers is significantly smaller than the evaluation data (505 papers) we used in the main body of this paper. Despite not designing a task specific method, we place second on the leaderboard, showing the powerful capabilities of LLMs in long document generation.
Appendix E Example Outputs
We provide examples of the outputs generated with and without the intermediate representation below. The documents in all examples are generated with GPT4 (gpt4-32k). Figure 5 includes example slide generations, Figure 6 includes example blog generations, and Figure 7 includes example poster generations.
(a) Document generated without intermediate representation. This example is not cropped.
(b) Document generated with intermediate representation. This example is cropped for space and includes an additional 4 slides that are not included for space.
Figure 5: The above documents are example slides generated by GPT4 (gpt4-32k) with and without the intermediate representation. We can see that without the intermediate step, the model did not generate a true slide deck.
(a) Document generated without the intermediate representation. This example is not cropped.
(b) Document generated with the intermediate representation. This example is cropped for space and includes an additional page of text.
Figure 6: The above documents are example blog posts generated by GPT4 (gpt4-32k) with and without the intermediate representation. We can see that without the intermediate representation, the model did not properly format the LaTeX file for compilation.
(a) Document generated without the intermediate representation. This example is cropped for space and includes an additional 3 slides.
(b) Document generated with the intermediate representation. This example is cropped for space and includes an additional 4 slides.
Figure 7: The above documents are example posters generated by GPT4 (gpt4-32k) with and without the intermediate representation. We found that GPT4 often generates slide decks in place of posters. We can see that the document generated without the intermediate representation contains more verbose panels and includes less formatting.